OpenStudy (anonymous):

integration question how would I find indefinate intergal (tanx)^5(secx)^3 dx

OpenStudy (anonymous):

calculator

OpenStudy (anonymous):

jk do some tricks with your trig knowledge

OpenStudy (anonymous):

what trig knowledge?

OpenStudy (anonymous):

lol I'll give it a try --I'm assuming you mean trig identities?

OpenStudy (anonymous):

yeah, you know tanx =sinx/cosx you also know secx=1/cos

OpenStudy (anonymous):

you also know that d/dx(tanx)=sec^2(x)

OpenStudy (anonymous):

I do now!

OpenStudy (luigi0210):

You should use \( \frac{d}{dx} (secx)=tanx~secx \)

OpenStudy (anonymous):

derivative of secx=tanxsecx

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sin^2(x)=1-cos^2(x) I solved it this way

OpenStudy (anonymous):

\[\int\limits \tan^5(x)\sec^3(x) dx\] - make the powers even \[\int\limits \tan^4(x)\sec^2(x)\sec(x)\tan(x)dx\] let u=sec(x), therefore du=sec(x)tan(x)dx \[\int\limits (\tan^2(x))^2\sec^2(x)\sec(x)\tan(x)dx = \int\limits (\sec^2(x)-1)^2\sec^2(x)\sec(x)\tan(x)dx\] \[= \int\limits (u^2-1)^2u^2du = \int\limits (u^4-2u^2+1)u^2du = \int\limits (u^6-2u^4+u^2)du\] Integrate \[(u^7/7)-(2u^5/5)+(u^3/3)+C\] Replace u with sec(x), then you're done

OpenStudy (mathmate):

Another method: Convert everything to cos(x) and set u=cos(x) \(\rm \int tan^5(x)sec^3(x)dx\) \(\rm= \int sin^4(x)/cos^8(x)~~sin(x)dx\)... set u=cos(x), du=-sin(x)dx \(\rm= \int -(1-u^2)^2/u^8 du\) \(\rm= \int -(1/u^8-2/u^6+/u^4) du\) \(\rm= 1/(7u^7)-2/(5u^5)+1/(3u^3) + C\) .... here u=cos(x) as @ShaeWX had it.

OpenStudy (anonymous):

trig was never my strong suit. Thanks guys

OpenStudy (anonymous):

I'm going to keep working on it until is sinks in

OpenStudy (mathmate):

@pblaesre That's the best way! :)